OpenStudy (anonymous):

If f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

first find g(-4) g(-4) = 8- (-4) g(-4) = 12 Now plug g(-4) in f(x), you get f(g(-4)) = f(12) f(12) = 12^2 - 2(12) + 9 f(12) = 144 -24 + 9 f(12) = 129

OpenStudy (anonymous):

thanks much! ^